a body is so projected in the air that the horizontal range coveredd b he body is equal to the maximum verticle height attained by the body during the motion. Fing the angle of projection?..........sir plz give the reason and thanks alot in advance

Grade:11

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
hoz. range =(u^2*sin 2 theta)/g
max. vertical height= (u^2 * sin^2 theta)/2g
as both are equal, (u^2*sin 2 theta)/g = (u^2 * sin^2 theta)/2g
so,sin 2 theta =sin^2 theta /2
=>2cos theta=sin theta /2
=> tan theta=4
=> theta = tan^-1 (4) 
so angle of inclination is tan^-1 (4). [Ans]
Hope this helped.
Thanks,
Soumya Ranjan Mohanty.